# Exploring the AGN Fraction of a Sample of JWST’s Little Red Dots at : Overmassive Black Holes Are Strongly Favored

### v1

@@ Abstract
JWST is revolutionizing our view of the early Universe by pushing the boundaries of detectable galaxies and black holes in redshift (upward) and mass (downward). The Little Red Dots (LRDs), detected by several surveys at , present a significant interpretational challenge, as their spectral energy distributions can mimic both AGN and stellar population templates. This study analyzes eight LRDs from the CEERS survey, utilizing NIRCam and MIRI photometry. By performing SED fitting across a vast parameter space, we explore a broad range of AGN fractions, defined as the ratio of the monochromatic luminosities (AGN, galaxy, and dust) at 0.53 rest-frame. We find that the SEDs of all the LRDs investigated are consistent with having significant AGN contributions, with the best-fitting AGN fractions ranging between 40% and 85%. Moreover, assuming these LRDs do indeed host AGN, we can place limits on their black hole masses using the inferred AGN bolometric luminosities and adopting the Eddington limit. We find that, independent of the specific AGN fraction adopted, the LRDs’ black holes are significantly overmassive relative to their host galaxies—by dex, and up to dex in the most extreme cases—compared to the local relation. The presence of overmassive black holes in the high- Universe may provide the strongest evidence yet of heavy black hole seeding occurring during the cosmic dark ages.
@@ 1 Introduction
The prevalence of such overmassive black holes brings into question the seeding mechanism responsible for the formation of such high- systems (Pacucci et al., 2023; Volonteri et al., 2023; Natarajan et al., 2024; Pacucci & Loeb, 2024). Heavy seed scenarios (see, e.g., Volonteri et al. 2008, and the reviews by Woods et al. 2019; Inayoshi et al. 2020) were first introduced to explain the extremely massive SMBHs observed at (see, e.g., Mortlock et al. 2011; Wu et al. 2015; Bañados et al. 2018), which would otherwise lack the necessary time to growth from standard, low-mass seeds. These heavy seeding models refer to various channels that may have shaped the formation of black holes in the early Universe, such as direct collapse black holes (Loeb & Rasio, 1994; Bromm & Loeb, 2003; Lodato & Natarajan, 2006). Some recent studies (e.g., Pacucci et al. 2023; Natarajan et al. 2024; Scoggins et al. 2023; Pacucci & Loeb 2024) suggest that the high- overmassive black holes observed by JWST may be the direct result of this heavy seeding scenario.
@@ 2 Data
This study is based on the sample of EROs by Barro et al. (2024), which was the first study to analyze these remarkably red and compact sources thoroughly.
Barro et al. (2024) report the discovery of 37 EROs, whose selection is based on NIRCam colors F277W-F444W 1.5 mag. These objects are selected from the JWST’s Cosmic Evolution Early Release Science Survey (CEERS, Finkelstein et al. 2023), which surveys arcmin2 of the Extended Groth Strip (EGS). Out of the 37 sources in Barro et al. (2024), we select a sub-sample of eight sources, which are characterized by spectra acquired with NIRSpec, within and , and MIRI, within and .
These objects were observed as very red galaxies in the JWST/NIRCam filters and additionally detected in either the NIRSpec filters or the MIRI filters. A common feature in all eight galaxies is a bump in the redder part of the photometric spectrum, often attributed to AGN emission.
2.1 A Note on Terminology: Extremely Red Objects vs. Little Red Dots
The name “Extremely Red Objects”, or EROs, precedes that of “Little Red Dots”, or LRDs, which has now become a standard nomenclature for the compact and strikingly red objects commonly found by JWST between and (Kocevski et al., 2024).
The color selection performed by Barro et al. (2024) for the EROs is consistent with the current definition of LRDs (see, e.g., the extensive catalog provided in Kocevski et al. 2024); however, the selection used in Barro et al. (2024) was slightly different. This paper selected extremely red sources with F277W-F444W , but no selection on morphology or rest-frame UV color was performed. Remarkably, the vast majority of sources with the F277W-F444W color selection are also point-like and have blue F150W-F200W colors, which fit the current definition of LRDs (Labbe et al., 2023; Kokorev et al., 2023; Kocevski et al., 2023, 2024). However, the subset studied in this paper do not have rest frame optical spectroscopy; most objects in the general population of LRDs have broad line emission consistent with Type 1 (unobscured) AGN (Greene et al., 2024). Additionally, and as discussed below, the population studied here are best modeled as Type 2 (obscured) AGN, suggesting they may represent a particular subset of LRDs.
@@ 3 Methods: SED modeling
Star Formation History | e-folding time (main) | 0.1, 0.4, 0.8, 1, 2, 3, 4, 5 Gyrs | |
Main stellar age | 0.1, 0.2, 0.4, 0.5, 0.6, 0.7 Gyrs | ||
e-folding time (burst) | 0.01, 0.05, 0.1 Gyrs | ||
sfhdelayed | Age of the burst population | 0.01, 0.03, 0.05 Gyrs | |
Burst fraction | 0, 0.001, 0.01, 0.1, 0.2, 0.3, 0.4, 0.5 Gyrs | ||
@@
AGN emission | Optical depth at 9.7 | 7 | |
Opening angle | 40° | ||
@@
Skirtor 2016 | AGN contribution | fracAGN | 0.1 to 0.99 |
Wavelength range for fracAGN | 0.53 m | ||
Extinction in polar direction | E(B-V) | 1, 2, 3, 4, 5, 6 |
@@
We fit the available photometry of the eight LRDs using SEDs built on a combination of two emission templates: stellar and AGN. We selected templates from the ones provided in CIGALE to vary only a small subset of parameters to focus our analysis on how the AGN fraction affects the SED fitting results. Table 1 lists the selected CIGALE templates and the parameters that were kept free, alongside the parameters where we adopted a different value from the default value. We note that we adopted the fiducial values of redshift for each source, as provided by Barro et al. (2024); hence, CIGALE does not fit redshifts.
@@
For the star formation history, we use a standard delayed model with exponential burst: sfhdelayed. We allow the -folding time to vary from 0.1 to 5 Gyrs and the main stellar age from 0.1 Gyrs to the age of the Universe at the observed . These values are typically used in the literature for similar analyses. As these galaxies could be undergoing rapid evolution in their stellar population, we allow a range of the e-folding time for the burst population (0.01 to 0.1 Gyr), the age of the burst population (0.001, 0.05 Gyr), and the burst fraction (0 to 0.5).
@@
For the simple stellar population, we use the bc03 module (Bruzual & Charlot, 2003), with a Chabrier (2003) IMF. Although the module defaults to solar metallicity ( = 0.02), we find that changing this parameter had little effect on our modeling. We included nebular emissions using the nebular module Bruzual & Charlot (2003). For dust attenuation, we employ the dustatt_modified_starburst module based on Charlot & Fall (2000). We employ the dl2014 module Draine et al. (2014) for galactic dust emission. For spatially unresolved sources at high redshift, there is little information regarding the relative contributions of the torus and the interstellar medium to the total dust attenuation (Hickox & Alexander, 2018). For our analysis, we do not vary dust attenuation on a galactic scale, instead choosing to focus on the torus for simplicity.
For the AGN component, we employ the skirtor2016 module based on clumpy AGN models from Stalevski et al. (2012) and Stalevski et al. (2016). We fix the parameter (i.e., the AGN optical depth at 9.7 m) to 7 and the opening angle of the torus to 40°. As skirtor uses the inclination to determine the AGN type, we allow this parameter to vary between 30°(type II) and 70°(type I)111 Note that CIGALE uses a convention where type II AGN is defined as [0°, 90°-], and type I AGN is defined as [90°-, 90°], where is the opening angle. This definition is inverted with respect to normal conventions.
@@
The AGN fraction in skirtor is defined as the ratio between the AGN luminosity and the total luminosity over a wavelength range defined by the user. We set the (rest-frame) wavelength to 0.53 m for our analysis. We then fix the AGN fraction between 5% and 99% to determine the best set of parameters at a given AGN fraction. We changed the torus dust attenuation parameter from its default value of 0.003 and allowed it to vary between 1 and 6, as we expect this range of E(B-V) values to better match the level of obscuration reported for JWST galaxies at .
@@ 4 Results
4.1 SED models
During the SED modeling process, we found specific phenomenological trends that persisted in all analyzed LRDs. The residuals of the CIGALE models, as in the example of Fig. 1, are typically noisier at longer wavelengths. We highlight two possible explanations for this effect. First, the nebular emissions in CIGALE seem not to be well accounted for in the composite spectrum. Despite our efforts, the nebular peaks in Fig. 1 (at ) consistently show the largest residuals from the model. Second, the longer IR wavelengths are covered by MIRI, which has the shallowest coverage for the set of photometry we are modeling. In addition to larger residuals, we find that, for the LRDs that lack MIRI coverage, CIGALE modeling is biased towards an IR luminous AGN template. This strongly suggests that deep MIRI coverage is needed to accurately constrain the AGN contribution in the SEDs of LRDs (Greene et al., 2024).
Our initial approach was to test the notion that LRDs are high- obscured AGNs by fitting our sample objects as type I (unobscured) and type II (dust-obscured) AGNs. However, we found that type I AGN models failed to produce the correct emission profile to match the IR bump in our SEDs, resulting in significantly worse reduced when compared to type II models. Therefore, we decided to fit all our LRDs as Type II AGNs.
@@ 4.2 Stellar-Dominated or AGN-Dominated?
This analysis aims to constrain the nature of the LRDs while accounting for our lack of information on precisely how much the AGN contributes to the observed SED. To this end, we perform a complete analysis of how a varying AGN fraction modifies the location of the galaxy in the plane; we then use the reduced statistic provided by CIGALE to determine the best-fit parameters.
@@
This process provided estimates for the black hole and stellar masses for any given AGN fraction. An example of this study of the parameter space is shown in Figure 2; as the AGN fraction increases (i.e., higher AGN contribution to the overall SED), so does the inferred black hole mass, which then corresponds to a decrease in stellar mass. This observed behavior matches our expectations based on how CIGALE uses composite models to construct SEDs.
The sharp drop that can be seen at an AGN fraction of is likely due to the absence of MIRI coverage at longer wavelengths, leading to a preference for a highly reddened AGN component. By considering models where the AGN is highly reddened, CIGALE no longer varies the AGN contribution in response to a change in AGN fraction. Rather, the flux of the galaxy component is lowered to compensate for the additional contribution to the optical luminosity from the IR-luminous AGN component.
For the eight sources analyzed, the AGN fractions yielding the lowest chi-squared values range between 40% and 85%, suggesting that the SEDs of these systems are consistent with a strong AGN contribution. However, the precise value of the AGN fraction is not well constrained; further analysis, such as spectroscopic follow-up, would be required to confirm the contribution of the AGN to the total luminosity in these sources.
@@ 4.3 Location in the Plane
The position of the eight LRDs on the plane for their best-fit models are shown by stars in Fig. 3. Generally, we find black hole masses ranging from to .
We have also plotted the Reines & Volonteri (2015) relation and the inferred high relation from Pacucci et al. (2023). Except for NC5-9553, we find that the black holes in these galaxies are overmassive by orders of magnitude (and up to in extreme cases) compared to their local Universe counterparts. Hence, these specific LRDs are well described by a scaling relation in the plane that favors overmassive black holes, such as the one presented in Pacucci et al. (2023).
@@
The same figure displays the track in the plane that is due to a varying AGN fraction (in the range 5% to 99%) for each LRD. The tracks show that even assuming a modest AGN contribution ( 10%), SED fitting strongly suggests the presence of an overmassive black hole. Furthermore, for all eight LRDs, the tracks evolve to the plot’s left side, showing that as the AGN fraction increases, the objects become more overmassive on the plane. We argue that this result shows that if the LRDs observed by JWST are AGN hosts, they must be overmassive regardless of how much the AGN contributes to the overall SED.

### v2

@@ Abstract
JWST is revolutionizing our view of the early Universe by pushing the boundaries of detectable galaxies and black holes in redshift (upward) and mass (downward). The Little Red Dots (LRDs), detected by several surveys at , present a significant interpretational challenge, as their Spectral Energy Distributions (SED) can mimic both AGN and stellar population templates. This study analyzes 19 LRDs from the JADES survey, utilizing NIRCam and MIRI photometry. By performing SED fitting across a vast parameter space, we explore a broad range of AGN fractions, defined as the ratio of the monochromatic luminosities (AGN, galaxy, and dust) over a specified wavelength range, 0.4 - 0.7 rest-frame. We find that 17 of the 19 LRDs investigated are consistent with having significant AGN contributions, with best-fitting AGN fractions ranging between 20% and 70%, while one galaxy shows a low AGN contribution (2%) and another appears to be purely star-forming. Moreover, assuming these LRDs do indeed host AGN, we can place limits on their black hole masses using the inferred AGN bolometric luminosities and adopting the Eddington limit. We find that, independent of the specific AGN fraction adopted, the LRDs’ black holes are significantly overmassive relative to their host galaxies—by dex, and up to dex in the most extreme cases—compared to the local relation. The presence of overmassive black holes in the high- Universe may provide the strongest evidence yet of heavy black hole seeding occurring during the cosmic dark ages.
@@ 1 Introduction
The prevalence of such overmassive black holes brings into question the seeding mechanism responsible for the formation of such high- systems (Pacucci et al., 2023; Volonteri et al., 2023; Natarajan et al., 2024; Pacucci & Loeb, 2024). Heavy seed scenarios (see, e.g., Volonteri et al. 2008, and the reviews by Woods et al. 2019; Inayoshi et al. 2020) were first introduced to explain the extremely massive SMBHs observed at (see, e.g., Mortlock et al. 2011; Wu et al. 2015; Bañados et al. 2018), which would otherwise lack the necessary time to grow from standard, low-mass seeds (Pacucci & Loeb, 2022). These heavy seeding models refer to various channels that may have shaped the formation of black holes in the early Universe, such as direct collapse black holes (Loeb & Rasio, 1994; Bromm & Loeb, 2003; Lodato & Natarajan, 2006). Some recent studies (e.g., Pacucci et al. 2023; Natarajan et al. 2024; Scoggins et al. 2023; Pacucci & Loeb 2024) suggest that the high- overmassive black holes observed by JWST may be the direct result of this heavy seeding scenario.
@@ 2 Data
This study is based on the sample of LRDs by Pérez-González et al. (2024), which is currently the deepest publicly available catalog of galaxies observed with the Mid-Infrared instrument (MIRI) on JWST. As discussed in detail later on, this long-wavelength data is essential to constraining the presence of AGN in LRDs. Hence, we focus our efforts on the dataset containing the longest wavelength MIRI observation of LRDs.
Pérez-González et al. (2024) reports the discovery of 31 LRDs, whose selection is based on NIRcam colors F277W-F444W 1 mag, F150W-F200W 0.5 mag, and F444W 28 mag. These objects are selected from the JWST Advanced Deep Extragalactic Survey (JADES, Eisenstein et al. 2023), with matched Mid-IR observation from the SMILES survey. We select a sub-sample of 19 sources, which are characterized by MIRI detection, within and . We only adopt fluxes in the NIRCam filters (F090W, F115W, F150W, F182M, F200W, F210M, F277W, F335M, F356W, F410M, F430M, F444W, F460M, F480M), and MIRI filters (F560W, F770W, F1000W, F1280W, F1500W, F1800W, F2100W, F2550W). Similar to Pérez-González et al. 2024, we impose a 5 upper limit for filters where there are non-detections. We do not use any HST fluxes or upper limits in our analysis.
18 out of the 31 galaxies had spectroscopic redshifts from the survey where they were selected. Photometric redshifts were estimated for the remaining galaxies using SED fitting with EAZYPY. The total Pérez-González et al. (2024) sample had redshifts between z = 4 - 8.5, with our selected sub-sample spanning a slightly smaller redshift range of z = 4 - 8. 15 of the 18 LRDs with spectroscopic redshifts are included in the analysis presented here.
@@ 3 Methods: SED modeling
Star Formation History | e-folding time (main) | 0.1 - 5 (30 steps) Gyrs | |
sfhdelayed | Main stellar age | 0.1 - 2 (30 steps) Gyrs | |
@@
Lyman escape fraction | 0.09 | ||
Dust attenuation | Color excess of nebular lines | E(B-V)line | 0.1 |
AGN emission | Opening angle | 40° | |
@@
Skirtor 2016 | AGN contribution | fracAGN | 0.02 - 0.2 (step 0.02), 0.2 - 0.9 (step 0.1), 0.99 |
Wavelength range for fracAGN | 0.4 - 0.7 m | ||
Extinction in polar direction | E(B-V) | 1 - 4 (16 steps) |
@@
We fit the available photometry of the eight LRDs using SEDs built on a combination of two emission templates: stellar and AGN. We selected templates from the ones provided in CIGALE to vary only a small subset of parameters to focus our analysis on how the AGN fraction affects the SED fitting results. Table 1 lists the selected CIGALE templates and the parameters that were kept free, alongside the parameters where we adopted a different value from the default value. We note that we adopted the fiducial values of redshift for each source, as provided by Pérez-González et al. (2024); hence, CIGALE does not fit redshifts.
@@
For the star formation history, we use a standard delayed model with exponential burst: sfhdelayed. We allow the -folding time to vary from 0.1 to 5 Gyrs and the main stellar age from 0.1 Gyrs to the age of the Universe at the observed . These values are typically used in the literature for similar analyses. For the simple stellar population, we use the bc03 module (Bruzual & Charlot, 2003), with a Chabrier (2003) IMF. Although the module defaults to solar metallicity ( = 0.02), we find that changing this parameter had little effect on our modeling. To model for nebular emissions from the HII regions, we use the nebular module from Inoue (2011). We model the physics of the Lyman alpha emission using results from Lin et al. 2024, which found a Lyman alpha escape fraction of 9% from H emitting galaxies observed by JWST between z = 4.9 - 6.3.
@@
For dust attenuation, we employ the dustatt_modified_starburst module based on the Calzetti et al. 2000 attenuation law, extended to lower wavelengths with Leitherer et al. 2002. The Ly result from Lin et al. 2024 employed a Small Magellanic Cloud dust law, yielding an E(B-V) color excess of 0.1, which we also adopt for our analysis. We employ the dl2014 module Draine et al. (2014) for galactic dust emission. For spatially unresolved sources at high redshift, there is little information regarding the relative contributions of the torus and the interstellar medium to the total dust attenuation (Hickox & Alexander, 2018). For our analysis, we do not vary dust attenuation on a galactic scale; instead, we choose to focus on the torus for simplicity.
For the AGN component, we employ the skirtor2016 module based on clumpy AGN models from Stalevski et al. (2012) and Stalevski et al. (2016). skirtor2016 uses three parameters, , , and to model the dust dynamics. The parameter sets the average edge on optical depth at 9.7 m (possible values: 3,5,7,9,11), is the power-law exponent that sets the radial gradient of the dust density (possible values: 0,0.5,1,1.5), and denotes the dust density gradient with polar angle (possible values: 0,0.5,1,1.5). Varying these parameters had little effect on the fit results but significantly increased the computation time for CIGALE. To reduce the compute time required, we fixed all three parameters. We adopt a modest opening angle of the torus at 40°. In skirtor, the inclination angle determines the AGN type; therefore, we fix this parameter to either 30°for Type I templates or 70°for Type II templates. We do not vary the inclination for the two AGN types because the fit results are generally insensitive to different viewing angles.
@@
The AGN fraction in skirtor is defined as the ratio between the AGN luminosity and the total luminosity over a wavelength range defined by the user. We set the (rest-frame) wavelength range to 0.4 - 0.7 m, to ensure that the AGN fraction is calculated over a range of wavelengths that is observed by JWST. We then fix the AGN fraction between 2% and 99% to determine the best set of parameters at a given AGN fraction. For the torus dust attenuation, we adopt a Calzetti et al. 2000 attenuation law, and we allow the torus reddening parameter to vary between 1 and 4, corresponding to an value of 3.63 and 14.52 respectively, as we expect this range of reddening values to better match the level of obscuration reported for JWST galaxies at . See Section 4 for a discussion on how fixing this parameter affects our SED fitting results.
@@ 4 Results
4.1 AGN nature of LRD
This analysis aims to constrain the nature of the LRDs while accounting for our lack of information on precisely how much the AGN contributes to the observed SED. To this end, we model our photometry using both Type I unobscured AGN and Type II obscured AGN templates. For all 19 galaxies in our analysis, there is a clear preference for the Type I AGN template, as it consistently returned better reduced s than the obscured Type II template. The Type II templates typically yielded reduced values an order of magnitude larger than the Type I fits, often with such poor statistics that CIGALE could not return a best-fit model. Therefore, the remainder of this paper will focus on the results of the SED fitting using the Type I unobscured AGN template.
In addition to this, the reddening levels that are recovered from the Type I templates are more in agreement with the current landscape of LRD, where a significant amount is observed to have broad line emission (Greene et al., 2024).
@@ 4.2 Stellar-Dominated or AGN-Dominated?
To constrain how the modeling of AGN contribution to the LRD SED affects the recovered stellar parameters, we perform a complete analysis of how a varying AGN fraction modifies the location of the galaxy in the plane; we then use the reduced statistic provided by CIGALE to determine the best-fit parameters.
@@
This process provided estimates for the black hole and stellar masses for any given AGN fraction. An example of the parameter space for one of our LRDs, JADES-79803, is shown in Figure 2; as the AGN fraction increases (i.e., higher AGN contribution to the overall SED), so does the inferred black hole mass, which then corresponds to a decrease in stellar mass. This observed behavior matches our expectations based on how CIGALE uses composite models to construct SEDs.
Of the nineteen sources analyzed, we found one galaxy, JADES-187025, likely to be a star-forming galaxy with no AGN component. 11 out of the nineteen galaxies are best fit by an AGN fraction that is less than 40% while the rest of the population had best-fit AGN fractions between 40% and 70%, suggesting that, in general, the SEDs of these systems are consistent with some, and in certain cases, strong AGN contribution. However, the precise value of the AGN fraction is not well constrained; further analysis, such as spectroscopic follow-up, would be required to confirm the contribution of the AGN to the total luminosity in these sources. Nevertheless, our results clearly indicate that AGN contribution is required to explain most LRDs in our sample, and that significant AGN contribution is required to explain the overmassive nature of the black holes hosted by LRDs.
@@ 4.3 Location in the Plane
The position of the 19 LRDs on the plane for their best-fit models are shown by squares in Fig. 3a. Generally, we find black hole masses ranging from to .
We have also plotted the Reines & Volonteri (2015) relation and the inferred high relation from Pacucci et al. (2023).
We find that for galaxies where the best fit AGN contribution is below 20%, the estimated black hole masses are in agreement with the local Reines & Volonteri 2015 relation. For galaxies where the best fit AGN contribution is moderate (20% fracAGN 40%), the black hole masses are overmassive by orders of magnitude compared to their local universe counterparts. For galaxies best fit by a significant AGN contribution (fracAGN 40%), the estimated black hole masses are overmasive by orders of magnitude (and up to in extreme cases). Hence, these specific LRDs with significant AGN contributions are best described by a scaling relation that favors overmassive black holes, such as the one described in Pacucci et al. (2023).
@@
The same figure displays the mean track111We use the mean tracks here to illustrate the general behavior of the plane, but it is not indicative of the true behavior. For the actual evolution of the plane, we direct the reader to fig. 5 in the plane that is due to a varying AGN fraction (in the range 2% to 99%) for each LRD. Interestingly, the tracks show that for an assumed modest AGN contribution ( 20%), the recovered AGN luminosity, and hence black hole masses are in agreement with the local Reines & Volonteri 2015 relationship. However, assuming significant AGN contribution ( 40%), SED fitting strongly suggests the presence of an overmassive black hole. Furthermore, for all 19 LRDs, the tracks evolve to the plot’s left side, showing that as the AGN fraction increases, the objects become more overmassive on the plane. We argue that this result shows that if the LRDs observed by JWST are AGN hosts, they are overmassive only if there is significant AGN contribution to the overall SED.
Figure 3a shows two galaxies for which the best fit AGN fraction is negligible (fracAGN = 2%), those two galaxies are Jades-187025 and Jades-204851. Jades-204851 as mentioned in 2, is the only broad line detected galaxy in our sample, and so it is concerning that it is only best fit by an AGN contribution of 2%. The surprisingly low best fit AGN contribution was suggestive of the two galaxies being star forming galaxies, instead of high redshift AGN hosts. For JADES-187025, this was indeed the case, as the SED fitting results for the stellar model returns a better reduced compared to the composite (AGN + stellar) model, and so we conclude that it is a dusty star forming galaxy. However, when we repeat the same analysis for JADES-204851, we find that the reduced for the stellar model is worse compared to the composite model.
Interestingly, Pérez-González et al. 2024 also recovered a low AGN contribution of approximately 10% when using a composite template of an unobscured QSO and a stellar template to fit the photometry for JADES-208451. Closer inspection of the CIGALE stellar model spectrum for JADES-204851 suggests that CIGALE could be underestimating the dust emission, which if correctly modeled could potentially lead to a much better SED fit for a stellar model. CIGALE modeling, much like the majority of the publicly available SED fitting codes, is based on an energy balance principle, where the absorbed light in the UV is remitted by the dust in the IR, this approach completely omits potential contributions from a hot dust component and has been found to be less reliable when fitting star forming galaxies at high redshifts (z 4, Haskell et al. 2023). And so it is plausible that JADES-204851 could be a high redshift star forming galaxy, whose broad line emission is due to kinematics and not AGN activity (Baggen et al., 2024).
4.3.1 AGN contribution at the Balmer limit, H∞
The spectral slope break at rest H∞ discussed in Setton et al. 2024 implies that an AGN model is inconsistent with explaining the spectroscopy of LRD. Here we asses the strength of the AGN contribution at the Balmer limit to determine the consistency of the composite model presented here under consideration of the spectral slope break. To constrain the AGN contribution at the Balmer limit, we define a new AGN fraction at rest H∞ (3645 Å). We then define a threshold of 50% to represent where the AGN contribution would impact the spectral slope at rest H∞. Figure 3b shows the location along the where the AGN contribution at the Balmer limit exceed 50 percent. We find that for most of the galaxies analyzed here (16/19) an AGN contribution that exceed 50% at the Balmer limit is consistent with a ( 90%) mid-IR AGN fraction. Such a high mid-IR AGN fraction is well beyond the typical ranges we recover as best fit AGN fractions, even in cases where the black hole masses are overmassive by up to a factor of 4, see Figure 3a. Therefore, we conclude that for the population of LRD with spectral slope break at rest H∞, a significant mid-IR AGN ( 70%) component that can produce an overmassive black hole result would preserve the spectral feature at rest H∞.
Figure 4 shows the continuum from the CIGALE model for the seven galaxies with F277W - F444W 1.7. For three out of the seven galaxies, there is a prominent spectral break at the Balmer limit, similar to what was reported by Setton et al. 2024. All three galaxies feature a best fitting AGN fraction in the rest-frame mid-IR between 40 and 70 percent. This shows that a reddened AGN component can account for the extra red emission in the optical for LRD while maintaining the H∞ spectral break. We must emphasize however, that this result does not rule out other possible sources for the optical red emission that is observed in LRD.
4.3.2 Photometric selection of sources
The LRD selection criteria adopted by Pérez-González et al. 2024 is largely consistent with other works (see Greene et al. 2024; Kocevski et al. 2023; Barro et al. 2024 ) in terms of color and compactness. The only discernible difference is the F277W - F444W magnitude criteria, where Pérez-González et al. 2024 uses a threshold of 1, other works used a threshold of 1.7. To test how this criteria affects our analysis, we restrict our sample to only include galaxies where the F277W - F444W is greater than 1.7. Adopting this color criteria reduced our galaxy catalog from nineteen to seven. Figure 3c shows the parameter space for the seven reddest galaxy in our catalog. Despite the smaller sample size, we find that our results still holds – the mid-IR AGN contribution needed to produce overmassive black holes is still significant ( 40%), while the threshold for the AGN emission to affect the spectral slope at rest H∞ is typically much higher than the best fit mid-IR AGN fraction.
4.4 AGN reddening and the dependence of SED fitting on MIRI observations
Our analysis identified two key factors that influence the behavior of the parameter space: (1) the availability of MIRI photometry and (2) the modeling of polar reddening in the AGN torus.
4.4.1 MIRI photometry
One of the key results from the LRD analysis is that the characteristic “v” shape of the SEDs makes the SED fitting comparable for both reddened star-forming (SF) templates and composite SF+AGN templates (Barro et al., 2024). However, most LRDs analyzed by Barro et al. 2024 had little to no MIRI photometry, with the deepest mid-IR data only reaching 0.77 m (rest-frame). We believe this lack of mid-IR data plays a critical role in the ambiguity of LRD photometry. To better understand the impact of missing mid-IR photometry on the SED fitting of LRDs, we removed all MIRI data from our catalog and re-ran the analysis.
We found that the SED fitting results, as shown in Figure 2, are highly sensitive to mid-IR photometry. Without MIRI photometry, the fitting returned comparable statistics for both AGN and non-AGN templates. Furthermore, the absence of MIRI data led CIGALE to prefer high polar reddening, with the average AGN polar reddening being 2 and typical values ranging between 4 and 20. This result underscores the importance of MIRI observations, as such extreme reddening would conflict with the significant portion of the LRD population observed to exhibit broad-line emission (Greene et al., 2024).
We also found that the depth of the included MIRI photometry affects the precision of CIGALE estimates of stellar masses. Specifically, we find that as the AGN contribution becomes maximal, the precision of the CIGALE stellar mass estimate worsens. However, we find that the precision becomes artificially enhanced as the included MIRI wavelengths becomes shorter. This result suggests that SED fitting analysis of LRD where the only considered AGN fraction is maximal could have some under-estimated errors on their stellar mass estimates, especially in cases where the MIRI photometry is not included.
4.4.2 Polar reddening of AGN torus
In addition to mid-IR photometry, we also found that the recovered stellar mass and AGN luminosity were highly sensitive to how we modeled the polar reddening, . Here we discuss how the modeling choices for the polar reddening, such as the attenuation law, and the flexibility of the reddening parameter,E(B-V), affected our analysis results.
Recent spectroscopic analysis of LRD (Greene et al., 2024; Maiolino et al., 2023) have commonly modeled the AGN polar reddening using a Small Magellanic Cloud (SMC) dust law from Pei 1993. The attenuation law used in our analysis follows fro the Balmer decrement analysis of LRDs from Brooks et al. 2024, as mentioned in section 3. To compare the effects of attenuation law on the SED fitting results, we re-ran our analysis using a SMC dust law. Using the SMC dust law resulted in a ( 10%) higher AGN luminosity compared to using the Calzetti et al. 2000 dust law.
In addition to the choice of attenuation law, we also tested how the model choice for the reddening parameter would affect our analysis. We initially allowed CIGALE to sample values between 1 and 4, corresponding to an range of 3.63 to 12, assuming a (Calzetti et al., 2000) extinction law. This sampling range is consistent with recent Balmer decrement analyses, which indicated that the ratio in LRDs suggests AGN reddening with (Brooks et al., 2024).
However, we found that the recovered parameter space performed poorly, as shown in Figure 6, especially at lower AGN fractions, where the reddening was typically higher, ranging between 1.4 and 3. Despite this, for all 19 galaxies in our sample, the best-fit AGN fraction occurred consistently at an value of 1. Motivated by this behavior, we re-ran our analysis with a fixed value of 1. As shown in Figure 5, this fixed polar AGN reddening behaves more smoothly across the parameter space compared to the results in Figure 6.